How do you solve?
x=1/5 y
3x+2y=26
Answer by Alan3354(69443) 
 How do you solve?
x=1/5 y
3x+2y=26
--------------
There are a number of ways, but since x is already shown as a function of y, substitution is probably the simplest.
x=1/5 y
3x+2y=26
Sub for x in the 2nd eqn
3*(y/5) + 2y = 26
Multiply by 5
3y + 10y = 130
13y = 130
y = 10
--------
x = y/5
x = 2
Avoid using 1/5 y, it could mean either (1/5)y or 1/(5y). Use y/5 instead, or add parentheses. Parentheses are free.
|
|
|